Typical Electronic Energy gaps in molecules are about 1.0 eV. In terms of temperature this gap is closest to


a. 10^2 K


b. 10^3 K


c. 10^4 K


d. 10^5 K

Grade:12

1 Answers

Akshay Meena
10 years ago
If you convert eV ( electron Volt ) into K (Kelvin) . Basically you are converting Energy (Joule) into Temperature (Kelvin)
1eV = 1.6 * 10 -19 J
For Converting it into K , We need Boltzman Constant as per the equation Suggested :
Kelvin (K) = ( 1eV / kB )         [Where , kB = 1.38 *10 -23 J/K ]
Kelvin (K) = ( 1.6 * 10 -19 J  /  1.38 *10 -23 J/K ) = 11604 K  equivalent to  10^4 K
So, It is Equivalent to Option (C)

Thanks & Regards
Akshay Meena, 
askIITians Faculty
M.Tech, IIT Kharagpur